14. Find the x-intercepts.

y=x^2-4x+4

I don't understand how to solve this type of problem at all. I need this explained in detail, the book is skipping steps and I can't figure out how or why they came up with their answer. Thanks in advance.

Answer by stanbon(75887)

Find the x-intercepts.
y=x^2-4x+4
-----------
To find the x-intercepts, let y=0 and solve for x as follows:
x^2-4x+4=0
(x-2)(x-2)=0
x=2
So the x-intercept is the point (2,0)
Hope this helps.
Cheers,
Stan H.